INAF queue observing, 2017 Sep 27-28

Observer: A. Rossi, R. Carini
Telescope Operator: G. Bechetti
Support Astronomer: D. Thomson
Instrument: LUCI1, LUCI2

Summary

At the sunset the sky is cloudy and rainy.

Night duration (12\xB0 el twilight) 2:00am- 12:20pm UT = 10 h 20m

Observing time: * *1.66hrs (16%)

Weather time loss: * *h (81%)

technical time loss * *0.35h (3%)

Data Summary

Project_RUN Objects Instrument Filter ExpTime Notes
dark   LUCI1,2      
Antoniucci HIP 116229 LUCI1,2     telluric for V733 useless

Details

*darks Piconcelli A11-A12 telluric Dit xndit 5x6, read-out lir.
  • luci1.20170928.0003-0007
  • luci2.20170928.0003-0007

*darks Rossi Dit xndit 20x3, read-out lir.
  • luci1.20170928.0009-0013
  • luci2.20170928.0013-0017

*darks Piconcelli A11 Dit xndit 180x1, read-out mer.
  • luci1.20170928.0014-0018
  • luci2.20170928.0018-0022

*darks Piconcelli A12 Dit xndit 240x1, read-out mer.
  • luci1.20170928.0019-0023
  • luci2.20170928.0023-0027

19:07 rain stop, but clouds at horizon. We wait for them to move away before opening the dome

*darks Antoniucci PVCep Dit xndit 60x1, read-out mer.
  • luci1.20170928.0024-0028
  • luci2.20170928.0028-0032

*darks Antoniucci PVCep Dit xndit 300x1, read-out mer.
  • luci1.20170928.0029-0033
  • luci2.20170928.0033-0037

19:59 we open the dome

03:30 we slew to telluric Hip 116229 (from V733) for PVCep (its telluric is too bright K~6), Antoniucci 8 program
  • luci2 error just after preset
  • luci1.20170928.0034-0035 useless

03:33 clouds again TO is checking if to close 04:20 misunderstanding , we lost 20 min because we thought we were closed for clouds.

04:20 we start again telluric, seeing 1.1"
  • acq luci1.20170928.0036-0038
  • acq luci2.20170928.0038-0040
  • 04:35 seeinf about 0.8''-0.9''
  • 04:50 error due to unstable wavelength center, lost 10 min
  • sci luci1.20170928.0039-0046 zJ - HK
  • sci luci2.20170928.0041-0049 zJ - HK check one position too much due to above error

04:53 slew to V733 Cep
  • acq luci1.20170928.0047-0049
  • acq luci2.20170928.0050-0052
  • 05:10 LUCI stuck!!!! Luci2 loses the guide stars for the clouds, it can not collimate so it can not take image, LUCI1 can guiding and goes on with the script.
  • 05:17 TO closes the Dome due to the clouds, Temperature and humidity more than limits

calibrations with LUCI
  • flats for Rossi luci1.0055-59, luci2.0058-0062

calibration for Antoniucci
  • flats zj ==luci1.0062-0066,luci2.0063-0067==without flexure compensation
  • arc zj Ne luci1.0067-0068, luci2.0069-0070 without flexure compensation
  • arc Ar luci1.0069-0070, luci2.0071-0072 without flexure compensation
  • flats HKspec ==luci1.0071-0075==without flexure compensation
  • arc Ne luci1.007-0079 without flexure compensation
  • arc Ar luci1.0078-0079 without flexure compensation
  • arc Xe luci1.0080-0081 without flexure compensation
  • it has not been possible to do the calibrations in LUCI2, the instrument stuck with this filter HKspec and the wavelength=1.93
Topic revision: r11 - 08 Nov 2017, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback